Question 9 Carla Company purchased machinery for $161,700 on January 1, 2017. It is estimated that the machinery will have a useful life of 20 years, salvage value of $14,700, production of 80,400 units, and working hours of 46,200. During 2017, the company uses the machinery for 21,714 hours, and the machinery produces 25,728 units. Compute depreciation under the straight-line, units-of-output, working hours, sum-of-the-years’-digits, and double-declining-balance methods. (Round intermediate calculations to 5 decimal places, e.g. 1.56487 and final answers to 0 decimal places, e.g. 5,125.) Depreciation Straight-line $ Units-of-output $ Working hours $ Sum-of-the-years’-digits $ Double-declining-balance
Answer #1

Calculate depreciation expense :

Straight line = (161700-14700/20) = 7350

Unit of output = (161700-14700/80400)*25728 = 47040

Working hours = (161700-14700/46200)*21714 = 69090

Sum of year digit = (161700-14700)*20/210 = 14000

Double decline = 161700*10% = 16170